OpenStudy (anonymous):

Which statement about y = x2 − x3 is true? It is not a function. It is an even function. It is neither an odd nor an even function. It is an odd function.

OpenStudy (anonymous):

for even functions: f(-x) = f(x) For odd functions: f(-x) = - f(x) for the given function f(-x) = (-x)^2-(-x)^3 = x^2+x^3 Therefore, It is neither an odd nor an even function.

OpenStudy (anonymous):

I just know its an even function

OpenStudy (anonymous):

ok Ima hold you to that

OpenStudy (anonymous):

you the best

OpenStudy (anonymous):

You are welcome :)